# Reproducing HotFlip for Corpus Poisoning Attacks in Dense Retrieval

### v1

@@ 1 Introduction
Finally, the adversarial passage generation pipeline of Zhong et al. [34] relies on prior knowledge of queries for which the attack is optimized (Figure 1, right). However, prior knowledge of user queries is not a realistic signal that we can rely on; especially in systems that value privacy. Our final research question (RQ3) is: How does HotFlip perform without prior query knowledge? We modify the method to use other passages instead of queries, and propose a new paradigm to measure the performance of HotFlip. We found that even corrupting as little as 0.01% of the corpus in this way can sometimes be devastating, depending on the target model and the size of the corpus.
@@ 5.1.2 Attack Efficiency Analysis
$\bullet$ Given that the MS MARCO training queries are more than three times greater than those of NQ, the Prep time for both pipelines on MS MARCO is correspondingly over three times longer than on NQ. However, the Search time for the same pipeline remains consistent across different datasets, indicating that it is determined by the retrieval model size rather than the dataset size.
@@ 5.3 RQ3: How does HotFlip perform without prior query knowledge?
In Section 3.2, the adversarial passage generation process relies on an unrealistic assumption that prior knowledge of queries is available. In practice, obtaining real user queries is challenging due to privacy constraints. Therefore, we design a new experimental setup where only part of the corpus is known, simulating the practical scenario of obtaining a corpus by manually entering human-generated queries and observing the ranking results. For our study, we utilize sampled batch data of the corpus, denoted as $\mathcal{C}_{b}$, to replace this manual process.
@@
Figure 4: The performance of query-agnostic attacks from HotFlip for RQ3. We attack ArguAna and FiQA with different retrievers, and the volume of injected adversarial passages $|\mathcal{A}|$ is decided by multiple percentage of corpus size.
@@ 5.4.1 $\ell$2 Norm of Embeddings
In the literature, bi-encoder retrievers are typically trained using the dot product, as it performs better when retrieving longer passages [26]. Consequently, all retrievers in this paper use the dot product as the similarity function: $\text{sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)\propto{\|E_{p}(p)\|}_{2}\rm{cos}\,\theta$, where $\rm{cos}\,\theta\in[0,1]$ and $\theta$ is the angle between the embeddings of the query and passage. We observe that in all of the above experiments, different retrieval models exhibited varying levels of robustness against HotFlip attacks. To investigate this, we analyze the $\ell$2 norm of the adversarial passages generated during the query-agnostic attacks (all percentages with all random seeds) on the ArguAna dataset in Section 5.3, as shown in Figure 5.

### v2

@@ 1 Introduction
Finally, the adversarial passage generation pipeline of Zhong et al. [34] relies on prior knowledge of queries for which the attack is optimized (Figure 1, right). However, prior knowledge of user queries is not a realistic signal that we can rely on; especially in systems that value privacy. Our final research question (RQ3) is: How does HotFlip perform without prior query knowledge? We modify the method to use other passages instead of queries, and propose a new paradigm to measure the performance of HotFlip. We found that even corrupting as little as 0.01% of the corpus in this way can sometimes decrease the effectiveness of the IR system, depending on the target model and the size of the corpus.
@@ 5.1.2 Attack Efficiency Analysis
$\bullet$ Given that there are three times more training queries in MS MARCO than in NQ, the Prep time for both pipelines on MS MARCO is correspondingly over three times longer than on NQ. However, the Search time for the same pipeline remains consistent across different datasets, indicating that it is determined by the retrieval model size rather than the dataset size.
@@ 5.3 RQ3: How does HotFlip perform without prior query knowledge?
In Section 3.2, the adversarial passage generation process relies on an unrealistic assumption that prior knowledge of queries is available. In practice, obtaining real user queries is challenging due to privacy constraints. Therefore, we design a new experimental setup where only part of the corpus is known, simulating the practical scenario of obtaining a corpus by manually entering human-generated queries and observing the ranking results. For our study, we utilize sampled batch data of the corpus, denoted as $\mathcal{C}_{b}$, to replace this manual process.2
@@
Figure 4: The performance of query-agnostic attacks from HotFlip for RQ3. We attack ArguAna and FiQA with different retrievers, and the volume of injected adversarial passages $|\mathcal{A}|$ is decided by multiple percentages of corpus size.
@@ 5.4.1 $\ell$2 Norm of Embeddings
In the literature, bi-encoder retrievers are typically trained using the dot product, as it performs better when retrieving longer passages [26]. Consequently, all retrievers in this paper use the dot product as the similarity function: $\text{sim}(q,p)=E_{q}(q)^{\top}E_{p}(p)\propto{\|E_{p}(p)\|}_{2}\rm{cos}\,\theta$, where $\rm{cos}\,\theta\in[-1,1]$ and $\theta$ is the angle between the embeddings of the query and passage. We observe that in all of the above experiments, different retrieval models exhibited varying levels of robustness against HotFlip attacks. To investigate this, we analyze the $\ell$2 norm of the adversarial passages generated during the query-agnostic attacks (all percentages with all random seeds) on the ArguAna dataset in Section 5.3, as shown in Figure 5.